40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 1 of 335  Not logged in ELOG logo
New entries since:Wed Dec 31 16:00:00 1969
ID Date Author Type Category Subject
  16859   Mon May 16 19:14:17 2022 AnchalUpdateBHDCamera set on AS path and BHDBS output path

[Anchal, Paco, Yuta]

Short summary:

  • We aligned AS path avoiding any clipping to the AP table where we setup a camera with a lens. To do this we had to move AS6 in -X direction for ~1cm.
  • We aligned the LO beam to be spatially matched on BHDBS with AS path.
  • The output of BHDBS was steered out and a GigE camera is setup to see this path. The camera is set to see the transmitted AS beam from BHD BS (and reflected LO beam).

Will add more details soon...

 

  16858   Mon May 16 16:13:01 2022 YehonathanUpdateBHDInitial BHD modeling: Damped suspension model

I was finally able to set up a stable suspension model with the help of Yuta and I'm now ready to start doing some MICH noise budgeting with BHD readout. (Tip: turns out that in the zpk function in Matlab you should multiply the poles and zeros by -2*pi to match the zpk TFs in Foton)

I copied all the filters from the suspension MEDM screens into a Matlab. Those filters were concatenated with a single pendulum suspension TF with poles at [0.05e-1+1i, 0.05e-1-1i] and a gain of 4 N/kg.

I multiplied the OLTF with the real gains at the DAC/DAC/OSEMs/Coil Driver and Coils. I ignore whitening/dewhitening for now. The OLTF was calculated with no additional ad-hoc gain.

Attachment 1 shows the calculated open-loop transfer function.

Attachment 2 shows OLTF of ETMY measured last week.

Attachment 3 shows the step and impulse responses of the closed-loop system.

 

 

Attachment 1: Damped_SUS.png
Damped_SUS.png
Attachment 2: ETMY_SUSPOS_GOL.pdf
ETMY_SUSPOS_GOL.pdf ETMY_SUSPOS_GOL.pdf
Attachment 3: SUS_Response.png
SUS_Response.png
  16857   Mon May 16 14:46:35 2022 TommyUpdateElectronicsRFSoC MTS Work

We followed the manual's guide for setting up MTS to sync on external signal. In the xrfdc package, we update the RFdc class to have RunMTS, SysRefEnable, and SysRefDisable functions as prescribed on page 180 of the manual. Then, we attempted to run the new functions in the notebook and read the DAC signal outputs on an oscilloscope. The DACs were not synced. We were also unable to get FIFOlatency readings. 

  16856   Mon May 16 13:22:59 2022 yutaUpdateBHDREFL and AS paths aligned at AP table

After Xarm and Yarm were aligned by Anchal et al, I aligned AS and REFL path in the AP table.
REFL path was alreasy almost perfectly aligned.

REFL path
 -REFL beam centered on the REFL camera
 -Aligned so that REFL55 and REFL33 RFPDs give maximum analog DC outputs when ITMY was misaligned to avoid MICH fringe
 -Aligned so that REFL11 give maximum C1:LSC-REFL11_I_ERR (analog DC output on REFL11 RFPD seemed to be not working)

AS path
 -AS beam centered on the AS camera. AS beam seems to be clipped at right side when you see at the viewport from -Y side.
 -Aligned so that AS55 give maximum C1:LSC-ASDC_OUT16 (analog DC output on AS55 RFPD seemed to be not working)
 -Aligned so that AS110 give maximum analog DC output

Attachment 1: REFLPOP.JPG
REFLPOP.JPG
Attachment 2: POPAS.JPG
POPAS.JPG
  16855   Mon May 16 12:59:27 2022 ChrisUpdateDAQDAQ troubleshooting

It looks like the RFM problem started a little after 2am on Saturday morning (attachment 1). It’s subsequent to what I did, but during a time of no apparent activity, either by me or others.

The pattern of errors on c1rfm (attachment 2) looks very much like this one previously reported by Gautam (errors on all IRFM0 ipcs). Maybe the fix described in Koji’s followup will work again (involving hard reboots).

Attachment 1: timeseries.png
timeseries.png
Attachment 2: err.png
err.png
  16854   Mon May 16 10:49:01 2022 AnchalUpdateDAQDAQ troubleshooting

[Anchal, Paco, JC]

Thanks Chris for the fix. We are able to access the testpoints now but we started facing another issue this morning, not sure how it is related to what you did.

  • The C1:LSC-TRX_OUT and C1:LSC-TRY_OUT channels are stuck to zero value.
  • These were the channels we used until last friday to align the interferometer.
  • These channels are routed through the c1rfm FE model (Reflected Memory model is the name, I think). These channels carry the IR transmission photodiode monitors at the two ends of the interferometer, where they are first logged into the local FEs as C1:SUS-ETMX_TRX and C1:SUS-ETMY_TRY .
  • These channels are then fed to C1:SCX-RFM_TRX -> C1:RFM_TRX -> C1:RFM-LSC_TRX -> C1:LSC-TRX and similar for Y side.
  • We are able to see channels in the end FE filtermodule testpoints (C1:SUS-ETMX_TRX_OUT & C1:SUS-ETMY_TRY_OUT)
  • However, we are unable to see the same signal in c1rfm filter module testpoints like C1:RFM_TRX_IN1, C1:RFM_TRY_IN1 etc
  • There is an IPC error shown in CDS FE status screen for c1rfm in c1sus. But we remember seeing this red for a long time and have been ignoring it so far as everything was working regardless.

The steps we have tried to fix this are:

  • Restart all the FE models in c1lsc, c1sus, and c1ioo (without restarting the computers themselves) , and then burt restore.
  • Restart all the FE models in c1iscex, and c1iscey (only c1iscey computer was restarted) , and then burt restore.

These above steps did not fix the issue. Since we have  the testpoints (C1:SUS-ETMX_TRX_OUT & C1:SUS-ETMY_TRY_OUT) for now to monitor the transmission levels, we are going ahead with our upgrade work without resovling this issue. Please let us know if you have any insights.

  16853   Sat May 14 08:36:03 2022 ChrisUpdateDAQDAQ troubleshooting

I heard a rumor about a DAQ problem at the 40m.

To investigate, I tried retrieving data from some channels under C1:SUS-AS1 on the c1sus2 front end. DQ channels worked fine, testpoint channels did not. This pointed to an issue involving the communication with awgtpman. However, AWG excitations did work. So the issue seemed to be specific to the communication between daqd and awgtpman.

daqd logs were complaining of an error in the tpRequest function: error code -3/couldn't create test point handle. (Confusingly, part of the error message was buffered somewhere, and would only print after a subsequent connection to daqd was made.) This message signifies some kind of failure in setting up the RPC connection to awgtpman. A further error string is available from the system to explain the cause of the failure, but daqd does not provide it. So we have to guess...

One of the reasons an RPC connection can fail is if the server name cannot be resolved. Indeed, address lookup for c1sus2 from fb1 was broken:

$ host c1sus2
Host c1sus2 not found: 3(NXDOMAIN)

In /etc/resolv.conf on fb1 there was the following line:

search martian.113.168.192.in-addr.arpa

Changing this to search martian got address lookup on fb1 working:

$ host c1sus2
c1sus2.martian has address 192.168.113.87

But testpoints still could not be retrieved from c1sus2, even after a daqd restart.

In /etc/hosts on fb1 I found the following:

192.168.113.92  c1sus2

Changing the hardcoded address to the value returned by the nameserver (192.168.113.87) fixed the problem.

It might be even better to remove the hardcoded addresses of front ends from the hosts file, letting DNS function as the sole source of truth. But a full system restart should be performed after such a change, to ensure nothing else is broken by it. I leave that for another time.

  16852   Fri May 13 18:42:13 2022 PacoUpdateAlignmentITMX and ITMY sat amp failures

[Yuta, Anchal, Paco]

As described briefly by JC, there were multiple failure modes going during this work segment. blush


ITMX SatAmp SAGA

Indeed, the 64 pin crimp cable from the gold sat amp box broke when work around ITMX chamber was ongoing. We found the right 64 pin head replacement around and moved on to fix the connector in-situ. After a first attempt, we suddenly lost all damping on vertex SUS (driven by these old sat amp electronics) because our c1susaux acromag chassis stopped working. After looking around the 1x5 rack electronics we noted that one of the +- 20 VDC Sorensens were at 11.6 VDC, drawing 6.7 A of current (nominally this supply draws over 5 Amps!) so we realized we had not connected the ITMX sat amp correctly, and the DC rail voltage drop busted the acromag power as well, tripping all the other watchdogs devil ...

We fixed this by first, unplugging the shorted cable from the rack (at which point the supply went back to 20 VDC, 4.7 A) and then carefully redoing the crimp connector. The second attempt was successful and we restored the c1susaux modbusIOC service (i.e. slow controls).


ITMY SatAmp SAGA

As we restored the slow controls, and damped most vertex suspensions, we noticed ITMY UL and SD osems were reading 0 counts both on the slow and fast ADCs. crying We suspected we had pulled some wires around when busy with the ITMX sat amp saga. We found that Side OSEM cLEMO cable was very loose on the whitening board. In fact, we have had no side osem signal on ITMY for some time. We fixed this. Nevertheless the UL channel remained silent... We then did the following tests:

  • Test PD mon outputs on the whitening card. We realized the whitening cards were mislabeled, with ITMX and ITMY flipped angry. We have labeled them appropriately.
  • Tested input DB15 cable with breakout board.
  • Went to the ITMY sat amp box and used the satellite box TESTER 2 on J1. It seemed correct.
  • We opened the chamber, tested the in-vacuum segments, they all were ok.
  • We flipped UR-UL OSEMs and found that the UL OSEM is healthy and works fine on UR channel.
  • We tested the in-air cable between satellite box and vacuum flange and it was ok too.
  • We suspected that the satellite box tester  is lying, so we replaced the satellite box with the spare old MC1 satellite box, and indeed that solved the issue.

DO NOT TRUST THE SATELLITE BOX TESTER 2.


Current state:

  • IMC locking normally.
  • All suspensions are damping properly.
  • Oplevs are not centered.
  • No flashing on either of the arms. We had no luck in ~20 min of attempt with just input injection changed.
  • On kicking PR3, we do see some flashing on XARM, which means XARM cavity atleast is somewhat aligned.
  • All remaining tasks before pumpdown are still remaining. We just lost the whole day.
  16851   Fri May 13 14:26:00 2022 JCUpdateAlignmentLO2 Beam

[Yehonathan, JC]

Yehonathan and I attempted to align the LO2 beam today through the BS chamber and ITMX Chamber. We found the LO2 beam was blocked by the POKM1 Mirror. During this attempt, I tapped TT2 with the Laser Card. This caused the mirror to shake and dampen into a new postion. Afterwards, when putting the door back on ITMX, one of the older cables were pulled and the insulation was torn. This caused some major issues and we have been able to regain either of the arms to their original standings.

  16850   Thu May 12 20:24:29 2022 yutaUpdateBHDPOX and POY investigation

[Anchal, Yuta]

We checked POX and POY RF signal chains for sanity check since Xarm cannot be locked in IR stably as opposed to Yarm.
POX beam seems to be healthy. This issue doesn't prevent us from closing the vacuum tank.

POY
 - RF PD has SPB-10.7+ and ZFL-500NL+ attached to the RF output.
 - At the demodulation electronics rack, SMA connectors are used everywhere.
 - With Yarm flashing at ~1, RF output has ~24 mVpp right after RF PD, ~580mVpp after SPB-10.7+ and ZFL-500NL+, and ~150mVpp at right before the demodulation box.
 - There is roughly a factor of 3 loss in the cabling from POY RF PD to the demodulation rack.
 - Laser power at POY RF PD was measured to be 16 uW

POX
 - RF PD doesn't have amplifiers attached.
 - At the demodulation electronics rack, N connector is used.
 - With Xarm flashing at ~1, RF output has ~30 mVpp right after RF PD, and ~20mVpp at right before the demodulation box.
 - Losses in the cabling from POX RF PD to the demodulation rack is small compared with that for POY.
 - Laser power at POX RF PD was measured to be 16 uW

Summary
 - POX and POY RF PDs are receiving almost the same mount of power
 - POY has larger error signal than POX because of RF amplifier, but the cable loss is high

Conclusion
 - There might be something in the electronics, but we can close the vacuum tanks

Attachment 1: POY.JPG
POY.JPG
  16849   Thu May 12 20:11:18 2022 AnchalUpdateBHDBHDBS Output beams steered out to ITMY table

I successfully steered out the two output beams from BHD BS to ITMY table today. This required significant changes on the table, but I was able to bring back the table to balance coarsely and then recover YARM flashing with fine tuning of ITMY.

  • The counterweights were kept at the North end of the table which was in way of one of the output beams of BHD.
  • So I saved the level meter positions in my head and removed those counterweights.
  • I also needed to remove the cable post for ITMY and SRM that was in the center of the table.
  • I installed a new cable post which is just for SRM and is behind AS2. ITMY's cable post is next to it on the other edge of the table. This is to ensure that BHD board can come in later without disturbing existing layout.
  • I got 3 Y1-45P and 1 Y1-0 mirror. The Y1-0 mirror was not installed on a mount, so I removed an older optic which was unlabeled and put this on it's mount.
  • Note that I noticed that some light (significant enough to be visible on my card) is leaking out of the 45P mirrors. We need to make sure we aren't loosing too much power due to this.
  • Both beams are steered through the center of the window, they are separating outside and not clipping on any of the existing optics outside. (See attachment 1, the red beam in the center is the ITMY oplev input beam and the two IR beams are the outputs from BHD BS).
  • Also note that I didn't find any LO beam while doing this work. I only used AS beam to align the path.
  • I centered the ITMY oplev at the end.

Next steps:

  • LO path needs to be tuned up and cleared off again. We need to match the beams on BHD BS as well.
  • Setup steering mirrors and photodiodes on the outside table on ITMY.
Attachment 1: signal-2022-05-12-201844.jpeg
signal-2022-05-12-201844.jpeg
  16848   Thu May 12 19:55:01 2022 TegaUpdateBHDAS path alignment

[Yuta, Tega]

We finally managed to steer the AS beam from ITMY chamber, through BS and IMC chambers, to the in-air AP table.

We moved the AS5 mirror north to its nominal position and we also moved the ASL lens on BS chamber back to its nominal position. Attached photos are taken after today's alignment work.

Attachment 1: AS.JPG
AS.JPG
Attachment 2: OMCchamber.JPG
OMCchamber.JPG
  16847   Thu May 12 17:20:08 2022 PacoUpdateAlignmentPOP Beam on CCD

[Paco]

Got POP beam centered on camera and nominally on the two PDs. Attachment #1 shows "carrier" camera.

Attachment 1: PXL_20220513_002047233.jpg
PXL_20220513_002047233.jpg
  16846   Thu May 12 13:46:59 2022 JcUpdateAlignmentPOP Beam

[Tega, JC]

Tega and I went in to adjust the POP being in the ITMX Table. The beam entered the table high, so we adjusted the this by adding mirrors (The highlighted in Turqoise are mirrors which adjust the pitch of the beam). All the mirrors are set and we are now in the process of adjusting the PD.

Attachment 1: IMG_0777.jpeg
IMG_0777.jpeg
  16845   Wed May 11 15:49:42 2022 JCUpdateOPLEV TablesGreen Beam OPLEV Alignment

[Paco, JC]

Paco and I began aligning the Green Beam in the BS Oplev Table. while aligning the GRN-TRX, the initial beam was entering the table a bit low. To fix this, Paco went into the chamber and correcting the pitch with the steering mirror. The GRN-TRX is now set up, both the PD and Camera. Paco is continuing to work on the GRN-TRY and will update later on today. 

In the morning, I will update this post with photos of the new arrangement of the BS OPLEV Table.


Update Wed May 11 16:54:49 2022

[Paco]

GRY is now better mode matched to the YARM and is on the edge of locking, but it more work is needed to improve the alignment. The key difference this time with respect to previous attempts was to scan the two lenses on translation stages along the green injection path. This improved the GTRY level by a factor of 2.5, and I know it can be further improved. Anyways, the locked HOMs are nicely centered on the GTRY PD, so we are likely done with the in-vac GTRY GTRX alignment.


Update Wed May 12 10:59:22 2022

[JC]

The GTRX PD is now set up and connected. The camera have been set to an angle because the cable to connect it is too thick for the camera to maintain its original position along the side. 

 

Attachment 1: IMG_0770.jpeg
IMG_0770.jpeg
  16844   Tue May 10 18:25:43 2022 PacoUpdateBHDGreen Transmission path

[Yuta, Paco]

We installed GRX_SM1, GRX_SM2, and finished aligning the GRY_SM1, and GRY_SM2 steering mirrors in the BS and IMC Chambers. GRY_SM1 was slightly misplaced (by ~ 2 inches), so we had to move it slightly. Luckily this didn't grossly misaligned the IMC, and we could recover quickly by touching MC1 & MC3 pitch, and MC1 slight yaw.

Then, Yuta installed GRX_SM1, and GRX_SM2 by repurposing two 45 AOI P-Pol GR transmission mirrors on the flowbench. Because one of the weights on the BSC was in the way of GRX_SM2, it was shifted it before installation. This probably shifted the balancing of the whole table. The GRY beam is still not lock-able to the YARM, so as a proxy for GRY transmission beam we used the slight GRX reflection from the BS, and noted slight clipping through PR3 (in transmission). This should probably be checked with GTRY.

We believe this is the last installation operation of this vent.


We made sure the WFS feedback loop is working, and realigned the arm cavities to be flashing.

  16843   Tue May 10 16:31:14 2022 TegaUpdateBHDITMX optlev return beam steered to QPD

Followed the steps below to complete the ITMX optlev installation. The ITMX optlev return beam now reaches its QPD without being blocked by the input steering mirror.

Although, I centered the ITMX optlev readout, this was not done when the XARM flashing is maximized bcos the IMC chamber was being worked on, so this should be done later when the IR beam is back.

Quote:

Next steps:

  • The Oplev beam paths need to be adjusted.
    • The ongoing beam steering mirror is blocking the returning beam, so the ongoing path needs to be changed.
    • First setup two irises to save ingoing path.
    • Then make space for the returning beam by changing the steering mirror positions.
    • Then recover the ingoing path to the center of irises.
    • Steer the returning beam to the QPD.
    • Then maximize the flashing on XARM and center the oplev to save this position.
  16842   Tue May 10 15:46:38 2022 JCUpdateBHDRelocate green TRX and TRY components from PSL table to BS table

[JC, Tega]

Tega and I cleaned up the BS OPLEV Table and took out a couple of mirrors and an extra PD. The PD which was removed is "IP-POS - X/Y  Reversed". In addition to this, the cable is zip-tied to the others located on the outside of the table in case this is required later on. 

Next, we placed the cameras and mirrors for the green beam into their postions. A beam splitter and 4 mirrors were relocated from PSL table and placed onto the BS Oplev table to complete this. I will upload the picture of the newly updated photo with arrows of the beam routes.

Attachment 1: IMG_0741.jpeg
IMG_0741.jpeg
Attachment 2: IMG_0753.jpeg
IMG_0753.jpeg
  16840   Mon May 9 23:18:44 2022 KojiUpdateBHDBHD Platform Progress status

I quickly ran the FEA model to check the resonant freqs of the BHD platform.
The boundary conditions were:

  • The platform was not loaded
  • FIxed constraints were given to the five legs

Don has optimized the cut-out size for the OMCs to increase the rigidity of the plate. Also, the ribbed grid is made at the bottom side.

The lowest mode is at 168Hz. Because there is no leg around, it seems reasonable to have this kind of mode as the fundamental mode.
The other mode lined up at 291Hz, 394Hz, 402Hz, ...
The mode freqs will be lower once the platform is loaded. But as the unloaded platform mode, these mode freqs sound pretty good numbers.

Attachment 1: Screenshot_2022-05-09_224145.png
Screenshot_2022-05-09_224145.png
Attachment 2: Screenshot_2022-05-09_224137.png
Screenshot_2022-05-09_224137.png
Attachment 3: bhd_platform_168.png
bhd_platform_168.png
Attachment 4: bhd_platform_291.png
bhd_platform_291.png
Attachment 5: bhd_platform_394.png
bhd_platform_394.png
Attachment 6: bhd_platform_402.png
bhd_platform_402.png
  16839   Mon May 9 22:19:06 2022 KojiUpdateBHDBHD Platform Progress status

[Don, Koji]

Don is working on finalizing the BHD Platform design. All the components on the BHD platform are almost populated and aligned.

Don is still working on the table legs so that we can detach the legs when we need to float the table in the future.
The BHD BS mount will have a third picomotor so that we can steer 3 dof with the mount while the remaining dof needs to be provided by the OMC.
The BHD BS position is going to be adjusted so that the incident and trans beams have sufficient clearance.
The OMC legs (kinematic mounts) need more work so that we can adjust their positions for initial setup while they can be the reference for the reproducible placement of the OMCs.
The OMCs are rigidly held with the legs. For the damping of the 1-kHz body bode, which has a relatively high Q, there will be a dissipative element touching the glass breadboard.

Attachment 1: Screenshot_2022-05-09_220220.png
Screenshot_2022-05-09_220220.png
Attachment 2: Screenshot_2022-05-09_222405.png
Screenshot_2022-05-09_222405.png
  16838   Mon May 9 18:49:05 2022 TegaUpdateBHDRelocate green TRX and TRY PDs/cams/optics from PSL table to BS table

[Paco, Tega]

Started work on the relocating the green transmission optics, cameras and PDs. Before removing the any of the optics, we checked and confirmed that the PDs and Cams are indeed connected to the GRN TRX/Y medm channels. Then added labels to the cables before moving them.

Plumbing:

  • Moved all power and signal cables for the PDs and cameras from PSL table to BS table. See attachment #1

Relocated Optics & PDs & Cameras:

  • TRX and TRY cameras
  • TRX and TRY PDs
  • 1 BS, 2 lens for PDs and a steering mirror, see Atachement #2

 

Attachment 1: IMG_20220509_184439943.jpg
IMG_20220509_184439943.jpg
Attachment 2: IMG_20220509_184154722.jpg
IMG_20220509_184154722.jpg
  16837   Mon May 9 18:43:03 2022 AnchalUpdateBHDITMX table layout corrected

As I went to correct the ITMX Oplev mirrors, I found that both mirrors were placed in very different positions than the design position. Part of the reason I think was to preserve outside oplev path, and party because a counterweight was in ITMXOL1 position. I had to do following steps to correct this:

  • I noted down level meter readings of the table before making any changes.
  • I removed the counter weight from near the center of the table.
  • I placed the Oplev mirrors in the nominal positions.
  • I placed the counter weight near previous position.
  • I moved a edge hanging counter weight to get back the level meter to its previous state coarsely.
  • Then I used dataviewer to find the previous OSEM PD monitor values and changed ITMX PIT and YAW to come closer to those PD values. And voila, I regained the flashing on Xarm. I nudged the ITMX pit and yaw bit more to maximize it.
  • I then went back to aligning the Oplevs properly.
  • Then I adjusted the POP mirrors to get the beam back through center of window. This was very tricky and took a lot of time.
  • Now the beam is going through near center and the oplev beams are far away enough from POP_SM5.
  • On the outside table, I noted the POP beam and the oplev beam. I corrected the pit of the returning beam to get the oplev beam at nominal height on outside table.

ITMX Sat Amp is flaky

[Anchal, Paco]

During the above work, i must have kicked the cable between the vacuum flange and the satellite amplifier box for ITMX. This disconnected all the OSEMs and Coils. We tried several things to debug this and finally found that nudging the connections on Sat Amp box brought the OSEMs and coils back online. Note that the connector was not partially out or in a state that obviously showed disconnection of the pins. I'm glad we are putting in new electronics soon for the vertex optics as well.


Next steps:

  • I showed Tega the returning oplev beam and the POP beam coming out of the ITMX chamber.
  • The Oplev beam paths need to be adjusted.
    • The ongoing beam steering mirror is blocking the returning beam, so the ongoing path needs to be changed.
    • First setup two irises to save ingoing path.
    • Then make space for the returning beam by changing the steering mirror positions.
    • Then recover the ingoing path to the center of irises.
    • Steer the returning beam to the QPD.
    • Then maximize the flashing on XARM and center the oplev to save this position.
  • POP beam needs to be directed to previous setup on far side of table.
    • The POP beam is coming out at the rising angle.
    • This is good for us if we do bit unconventional stuff and transfer the beam to other side of table at an elevated height. Given how close all the beams are coming out of the viewport, I think this is the best solution in terms of saving time.
    • Get the beam down to the old setup which was camera and photodiodes all aligned.

 

 

  16836   Mon May 9 15:32:14 2022 Ian MacMillanSummaryComputersQuantization Noise Calculation Summary

I made the first pass at a tool to measure the quantization noise of specific filters in the 40m system. The code for which can be found here. It takes the input to the filter bank and the filter coefficients for all of the filters in the filter bank. it then runs the input through all the filters and measures the quantization noise at each instance. It does this by subtracting the 64-bit output from the 32-bit output. Note: the actual system is 64 bit so I need to update it to subtract the 64-bit output from the 128-bit output using the long double format. This means that it must be run on a computer that supports the long double format. which I checked and Rossa does. The code outputs a number of plots that look like the one in Attachment 1. Koji suggested formatting a page for each of the filters that is automatically generated that shows the filter and the results as well as an SNR for the noise source. The code is formatted as a class so that it can be easily added to the IFOtest repo when it is ready.

I tracked down a filter that I thought may have lower thermal noise than the one that is currently used. The specifics of this will be in the DCC document version 2 that I am updating but a diagram of it is found in attachment 2. Preliminary calculations seemed to show that it had lower quantization noise than the current filter realization. I added this filter realization to the c code and ran a simple comparison between all of them. The results in Attachment 3 are not as good as I had hoped. The input was a two-toned sin wave. The low-level broadband signal between 10Hz and 4kHz is the quantization noise. The blue shows the current filter realization and there shows the generic and most basic direct form 2. The orange one is the new filter, which I personally call the Aircraft Biquad because I found it in this paper by the Hughes Aircraft Company. See fig 2 in paper. They call it the "modified canonic form realization" but there are about 20 filters in the paper that also share that name. in the DCC doc I have just given them numbers because it is easier. 

Whats next:

1) I need to make the review the qnoisetool code to make it compute the correct 64-bit noise. 

        a) I also want to add the new filter to the simulation to see how it does

2) Make the output into a summary page the way Koji suggested. 

3) complete the updated DCC document. I need to reconcile the differences between the calculation I made and the actual result of the simulation.

Attachment 1: SUS-ETMX_SUSYAW3_0.0.pdf
SUS-ETMX_SUSYAW3_0.0.pdf
Attachment 2: LowNoiseBiquad2.pdf
LowNoiseBiquad2.pdf
Attachment 3: quant_noise_floor.pdf
quant_noise_floor.pdf
  16835   Fri May 6 13:48:34 2022 AnchalUpdateBHDWFS loop instability fixed

[Yuta, Anchal]

We investigated why WFS loop wasn't working. It seemed like WFS1 PIT error signal has a huge offset which would push the loop to misalign all optics' PIT. So we did the following steps:

  • Cover the WFS with Aluminium foils. Run Sitemap>IOO>C1_IOO_WFS_MASTER>!Actions>Correct WFS DC offsets.
  • Then center the WFS beams on the QPDS while looking at C1:IOO-WFS1/2_PIT/YAW_DC channels.
  • Then Switch off WFS loop, Switch off Autolocker, toggler the PSL Shutter so that IMC unlocks and does not catch lock back, and tehn run Sitemap>IOO>C1_IOO_WFS_MASTER>!Actions>Correct WFS RF offsets.
  • The above script found significant changes required in WFS1 RF offsets. After this, we opened the shutter and WFS loops were working fine.
  16834   Thu May 5 17:50:48 2022 YehonathanUpdateBHDBHD Readout simulation

I have made a Simulink diagram to use in the MICH modeling (attachment) for the homodyne angle detection scheme. The model will be used for each optic separately and the noises will be combined in quadrature.

I gathered some more bits of info to fill the Simulink boxes. This is what I have so far:

Noise sources

# Displacement noises from gwinc
# OSEM sensing noise from the null stream
# OpLev noise from SUM channel + Seismic motion

freq = np.logspace(1, 4, 100)
coil_driver_noise = 1*freq/freq # pA/sqrt(Hz), elog 15846 
RIN = 1e-2*freq/freq #1/sqrt(Hz), elog 16082  
freq_noise = (1e6/freq**2) #Hz/sqrt(Hz), elog 15431
dark_noise = 1e-8 #V/sqrt(Hz) https://wiki-40m.ligo.caltech.edu/Electronics/RFPD/AS55
ADC_noise = 1e-6 #V/sqrt(Hz)
DAC_noise = 1e-6 #V/sqrt(Hz), elog 13003
 

TFs and gains

#POS->BHD from Finesse
#RIN->BHD from Finesse
#Frequency noise->BHD from finesse
#Control filters from MEDM
#Whitening filters from https://wiki-40m.ligo.caltech.edu/Electronics/WhiteningFilters
#Dewhitening filters from elog 12983  

DAC_gain = 6.285e-4 #V/cts, elog 16161

coil_driver_gain = 31 # elog 15534

coil_driver_TF = 0.016 #N/A per coil, elog 15846 
coil_R = 20e3 #Ohm,, elog 15846 
SUS_TF = 1/(0.25*freq**2) #m/N, single pendulum
OSEM_TF = 2*16384*1e3 #cts/m, https://wiki-40m.ligo.caltech.edu/Calibration
ADC_TF = 1638.4 #cts/V 
DCPD_responsivity = 0.8 #A/W
DCPD_transimpedance = 1e3 #V/A

Attachment 1: BHD_controls_40m_MICH.pdf
BHD_controls_40m_MICH.pdf
  16833   Thu May 5 17:05:31 2022 TegaUpdateBHDIMC & X/Y-arm alignment

[Yuta, Tega]

In order to setup POP camera and RFPD on the ITMX table, we decided to first work on the IMC and X/Y-arm alignment.

 

IMC alignment:

We zeroed IMC WFS outputs and aligned IMC manually to get IMC transmission of 1200 and reflection of 0.35.

 

Y-arm alignment:

We used the new video game tool that moves the pairs of mirrors - PR3 & ETMY, ITMY & ETMY - in common and differential modes. This brought the Y-arm flashing to 0.8. Note that we used the _OFFSET bias values for PR3 & ETMY alignment instead of the _COMM bias values.

 

X-arm alignment:

We repeated the same procedure of moving the pairs of mirrors - BS & ETMX, ITMX & ETMX - in common and differential modes but manually this time. This brought the X-arm flashing to ~1.0.

Attachment 1: IFO_aligment_Y_locking_on.png
IFO_aligment_Y_locking_on.png
  16832   Thu May 5 14:46:22 2022 AnchalUpdateBHDPOP beam height lowered, POP_SM4 raised

[Anchal, JC]

We first aligned the single arm cavity resonance for both arms to get maximum flashing. As we opened the chamber, I found that the POP beam was mostly hitting the POP_SM4 mirror but was clipping about 2 mm on the top edge.

I used TT2-PR3 to lower the injection beam angle and moved pairs of ITMY-ETMY, and ITMX-ETMX to recover as much flashing as I could in the both arms. Then, I moved PR2 in pitch from 49 to 71 to maximize the arm flashing again. After these steps, the POP beam was clearly within the POP_SM4 mirror but still in the upper half of the optic and there was maybe just a mm of clearance from the top edge. I decided to raise POP_SM4 mirror by 0.14" spacer. Now the beam is still in upper half of the mirror but has a good clearance from the edge.

The POP beam is coming outside in the in-air table at as a rising beam in the nominal path near the center of the window. This beam needs to be directed to the POP camera and RFPD on the far-side of the table.

Next steps:

  • In-air table work: Setup POP camera and RFPD.
  • In ITMX chamber, rotate ITMX Oplev mirror to clear the oplev beam off POP_SM5. Change oplev beam path outside accordingly.
  • Install green transmission from X-arm steering mirrors in BS chamber.
  • Install 4 steering mirrors in ITMY chamber at the two outputs of BHD BS to direct the beam outside.
  • Figure out POX11 rotation angle and get XARM locking as well.

 

  16831   Wed May 4 18:52:43 2022 TegaUpdateBHDPOP beam too high at POP SM4 on the ITMX chamber

[Yuta, Tega]

We needed to sort out the POXDC signal so we could work on X-arm alignment. Given that POXDC channel value was approx 6 compared to POYDC value of approx. 180, we decided to open the ITMX chamber to see if we could improve the situation. We worked on the alignment of POX beam but could not improve the DC level which suggests that this was already optimized for.  As an aside, we also noticed some stray IR beam from the BS chamber, just above the POX beam which we cold not identify.

Next we moved on to the POP beam alignment, where we noticed that the beam level on LO1 and POP_SM4 was a bit on the high side. Basically, the beam was completely missing the 1" POP_SM4 mirror and was close to the top edge of LO1. So we changed TT2 pitch value from 0.0143 to -0.2357 in order to move the beam position on POP_SM4 mirror. This changed the input alignment, so we compensated using PR2 (0.0 -> 49.0) and PR3 (-5976.560 -> -5689.800). This did not get back the alignment as anticipated, so we moved ITMY pitch from 0.9297 to 0.9107. All of these alignment changes moved the POP beam down by approx 1/5 of an inch from outside the mirro to the edge of POP_SM4 mirror, where about half of the beam is clipped.

 

Next Steps:

We need to repeat these aligment procedures with say 1.5 time the change in TT2 pitch to center the beam on POP_SM4 mirror.

Attachment 1: snnp_pr2_pr3_itmy_aligment.png
snnp_pr2_pr3_itmy_aligment.png
  16830   Wed May 4 15:08:59 2022 TegaUpdateBHDXARM alignment to get flashing

[Yuta, Tega]

We aligned the BS, ITMY, and ETMY PIT and YAW to get the flashing on X-arm whilst also keeping the flashing of Y-arm. From attachment 1, it is clear that POXDC photodiode is not receiveing any light, so our next task is to work on POX alignment.

 

Quote:
  • Align X-arm cavity and regain flashing.
  • Fix the Oplev path for ITMX.
  • Tune POX11 phase angle to get an error signal with which we can lock the cavity.
  • Finish AS beam path setup.

 

Attachment 1: IFO_XY_ARMS_Flashing.png
IFO_XY_ARMS_Flashing.png
  16829   Wed May 4 13:09:42 2022 AnchalUpdateBHDIMC table balanced again, IMC is locking, YARM is locking

[Paco, JC, Anchal]

We balanced the IMC table back again to point that got us 50% of nominal transmission from IMC. Then we tweaked the steering mirror for injection to IMC to get up to 90% of nominal transmission. Finally, we used WFS servo loop to get to the 100% nominal transmission from IMC. However, we found that the WFS loop has been compromised now. It eventually misaligns IMC if left running for a few minutes. This needs to be investigated and fixed.

Next steps:

  • Align X-arm cavity and regain flashing.
  • Fix the Oplev path for ITMX.
  • Tune POX11 phase angle to get an error signal with which we can lock the cavity.
  • Finish AS beam path setup.
  16828   Tue May 3 21:07:17 2022 YehonathanUpdateBHDBHD Readout simulation

I feel like there is an instability in my thought process on this. Before my tendency to try to scale and generalize this problem brings me to a full stop I will make small but quick progress.

First thing is to calculate the noise budget for a simple Michelson. The involved optics are:

  • ITMX
  • ITMY
  • BS
  • LO1
  • LO2
  • AS1
  • AS4

all sensed with OSEMs and OpLevs only.

Things to fetch:

1. OSEM sensing noise. Where do I get the null stream (AKA butterfly mode)?

2. Oplev noise, look at the SUM channel (or this elog)

3. Actuation TF. Latest elog.

4. Coil driver noise. Going to take the HP supply curve from this elog.

5. Seismic noise + Seismic stack TF. Or maybe just take displacement noise from gwinc.

6. Laser noise. Still need to search.

7. DAC noise. Still need to search.

 

 

Quote:

I'm planning on simulating the BHD readout noise in a manner very similar to the ALS noise model using Simulink. I've made a sketch of the model for the longitudinal DOFs (attached). A model for ASC will be similar but with more measurement devices (OpLevs, QPDs, WFSs).

I'm not pretending to simulate everything in this diagram on the first go, it is just a sketch of the big picture.

 

  16827   Tue May 3 21:05:23 2022 yutaUpdateBHDAS path alignment, removing a lot of green stuff

[Tega, Yehonathan, Koji, Yuta]

We tried to align AS path this afternoon.
IMC is not aligned now after the work today crying
Green mirrors/perisocope in IMC chamber were removed since some of them was clipping the AS beam, and this changed the balance of the IMC stack and thus MC1 and MC3 alignment.

Summary of changes:
ITMY chamber
  - Rotated AS2 in roll by 90 deg to have more aperture for the transmission (photo)
  - IR beams are now centered on AS1, AS2, AS3 and AS4 (photo, photo)

BS chamber
  - Moved ASL towards -X direction for about 1/4 inch
  - Installed GRY_SM2 at the nominal position (re-used GR_SM3 from IMC chamber)

IMC chamber
  - Removed green optics GR_SM4, GR_SM3, GR_PERI2L (GR_PERI2L is now stored at Xend)
  - Removed IFI camera mirrors FIV1, FIV2 (they are now stored at Xend) (photo, photo)
  - GR_SM4 mount is now reused as GRY_SM1 (Y2-2037-0 is now mounted instead of previously mounted Y2-LW1-2050-UV-45P/AR), and GRY_SM1 is installed at the nominal position (photo)
  - Moved weights to balance the stack

OMC chamber (we don't have OMC in this chamber...)
  - We swapped AS5 and AS6 so that the nobs comes in -X direction to have more spacing between AS beam and IMC REFL beam (photo)
  - Moved weights to balance the stack

What we did:
1. Misaligned ITMX and use ITMY reflected beam to align AS path
2. Centered the IR beam on AS1 using SR2
3. Centered the IR beam on AS2 and ASL using AS1. AS2 was rotated in roll by 90 deg to have more aperture for the transmisson. cool
4. Centered the IR beam on AS3 using AS2 nobs, centered the IR beam on AS4 by rotating AS3 in yaw.
5. "AS beam" (it turned out that what we are looking was actually not the AS beam!! Some stray light) was in +X direction by 1 inch or so at AS5. Moving AS5 to center the beam would clip IMC REFL beam. So we swapped AS5 and AS6 so that the nobs comes in -X direction to have more spacing between AS beam and IMC REFL beam.
6. Balanced OMC chamber stack again using IMC REFL beam as a referece (bring the IMC REFL beam to the reference red circle on the monitor).
7. Tweaked the alignment of TT1 and TT2 to have Yarm flashing to ~0.9 in TRY. angel
8. Moved AS5 towards +X by an inch or so to center the "AS beam."
9. Moved ASL towards -X direction for about 1/4 inch and re-centered the beam by AS1 to see if the "AS beam" gets far from IMC REFL at OMC chamber, but the "AS beam" didn't move much.
10. By blocking the beam from ITMY, we found that "AS beam" was not the actual one. frown
11. Opened IMC chamber and found that AS beam is blocked by the past optics.
12. Removed old green optics and IFI camera mirrors. GR_SM4 mount and GR_SM3 were reused as mentioned above.
13. Tried to balance IMC chamber stack to recover IMC alignment. We used IMC REFL beam as a reference, but it was hard to completely bring the IMC REFL beam to the reference red circle on the monitor. It is now off by a beam diameter or so. No IMC flashing now. crying

Next:
Theoretically, balancing IMC chamber stack would recover all the IFO alignment, but maybe tough. It is maybe easier to align MC1 and MC3 to have IMC locked. Assuming input pointing to IMC is not drifted too much, we should be able to recover Yarm flashing by tweaking TT1 alignment only. However, MC3 SD OSEM is at the edge of the range. We might have to balance the stack more or tweak SD OSEM position.

Attachment 1: log_AS2Rotated.jpg
log_AS2Rotated.jpg
Attachment 2: log_AS5andAS6swapped.JPG
log_AS5andAS6swapped.JPG
Attachment 3: log_IFIMirrorsBeforeRemoving.JPG
log_IFIMirrorsBeforeRemoving.JPG
Attachment 4: log_IFIMirrorsPlacedatXend.JPG
log_IFIMirrorsPlacedatXend.JPG
Attachment 5: log_NewGRY_SM1.JPG
log_NewGRY_SM1.JPG
Attachment 6: log_OnAS3.JPG
log_OnAS3.JPG
Attachment 7: log_OnAS4.JPG
log_OnAS4.JPG
  16826   Tue May 3 14:02:09 2022 AnchalSummaryBHDInstalled POP path in ITMX Chamber

[Anchal, JC]

I installed POP_SM4 and POP_SM5 in the ITMX chamber in the nominal positions. This must have affected the ITMX Oplev because I could see that one of the ITMX oplev beam was going through POP_SM5. It needs to be changed in order to follow the original plan. However, since POP_SM5 is a 1064 line mirror, it is transparent to the opleve beam, so maybe we can just use the ITMX oplev in the current fashion.

Next steps:

  • Get flashing back on the XARM.
  • Try to get the correct phase angle in POX11 so that we can lock XARM with IR too.
  • Inspect ITMX Oplev. The quadrant sum is low so maybe it needs adjustment in the in-ar table.
  • Check if ITMX oplev path needs to be adjusted inside the chamber.
  16825   Tue May 3 13:18:47 2022 JCUpdateVACRGA Pump Down

Jordan, Tega, JC

Issue has been resolved. Breaker on RP1 was tripped so the RP1 button was reporting ON, but was not actually on which continuously tripped the V6 interlock. Breaker was reset, RP1 and RP3 turned on. The V6 was opened to rough out the RGA volume. Once, pressure was at ~100mtorr, V4 was opened to pump the RGA with TP2. V6 was closed and RP1/3 were turned off.

RGA is pumping down and will take scans next week to determine if a bakeout is needed

Quote:

Jordan and I, in order to start pumpig down the RGA Volume, we began by opening V7 and VM. Afterwards, we started RP1 and RP3. After this, the pressure in the line between RP1, RP3, and V6 dropped to 3.4 mTorr. Next, we tried to open V6, although an error message popped up. We haven't been able to erase it since. But we were able to turn on TP2 with V4 closed. The pressure in that line is reporting 1.4 mTorr.

 

PRP on the sitemap is giving off an incorrect pressure for the line between RP1, RP3, and V6. This is verified by the pressure by the control screen and the physical controller as well. 

 

  16824   Mon May 2 17:51:30 2022 AnchalUpdateBHDPOX Beam aligned on RFPD, YARM locked

[Paco, Anchal]

We found that one of the Y1-1037-45P marked mirror that we used was actually curved. So we removed it and used a different Y1-1037-45P mirror, adjusted the position of the lens and got the beam to land on POX11 RFPD successfully.

Then in control room, we maximized the POX11_I_ERR PDH signal amplitude by changing C1:LSC-POX11_PHASE_R to 42.95 from -67.7. We kept the C1:LSC-POX11_PHASE_D same at 90. We were getting +/- 200 PDH signal on POX_I_ERR.

Then in our attempt to lock the XARM, when we ran the "Restore XARM (POX)" script, YARM locked!

We are not sure why the YARM locked, we might have gotten lucky today. So we ran ASS on YARM and got the transmission (TRY_OUT) stable at 1. The lock is very robust and retrievable.

Coming back to XARM, we realized that the transmission photodiode used for XARM was the low-gain QPD instead of the thorlabs high gain photodiode. The high-gain photodiode was outputing large negative counts for some reason. We went to the Xend to investigate and found that the high gain photodiode was disconnected for some reason. Does anyone know/remember why we disconnected this photodiode?

We connected the photodiode back and it seems to work normally. We changed the photodiode selection back to high gain photodiode for TRX and on 40 dB attenuation, we see flashing between 1.4 to 1.6. However, we were unable to lock the XARM. We tried changing the gain of the loop, played a little bit with the trigger levels etc but couldn't get it to lock. Next shift team, please try to lock XARM.

Quote:

[Paco, Anchal, Yuta]

We opened the BSC and ITMX chamber in the morning (Friday) to investigate POX11 beam clipping. We immediately found that the POX11 beam was clipping by the recently installed cable posts, so luckily no major realingment had to be done after reinstalling the cable post in a better location.


Because we had the BSC open, we decided to steer the AS1 mirror to align the AS path from ITMY all the way to the vertex chamber.  Relatively small AS1 offsets (of ~ 2000 counts each) were added on PIT / YAW to center the beam on ASL (there is slight clipping along PIT, potentially because of the AS2 aperture. We then opened the vertex chamber and located the AS beam with relative ease. We decided to work on this chamber, since major changes propagate heavily downstream (simply changing the IMC pointing).

Anchal removed old optics from the vertex chamber and we installed the steering pair of mirrors for AS path. This changed the balance of the vertex table by a lot. By using the MC REFL camera beam spot we managed to coarsely balance the counterweights and recover the nominal IMC injection pointing. Simply reenabling the IMC autolocker gave us high transmission (~ 970 counts out of the typical 1200 these days).

The final IMC alignment was done by Anchal with delicate PIT motion on the input injection IMC miror to maximize the transmission (to our satisfaction, Anchal's motion was fine enough to keep the IMC locked). The end result was quite satisfying, as we recovered ~ 1200 counts of MC transmission.

Finally, we looked at the arm cavity transmission to see if we were lucky enough to see flashing. After not seeing it, we adjusted TT1 / TT2 to correct for any MMTT1 pitch adjustment needed after the vertex table rebalancing. Suprisingly, we didn't take too long and recovered the nominal arm cavity pointing after a little adjustment. We stopped here, but now the vertex table layout is final, and AS beam still needs to be aligned to the vertex in-air table.

 

  16823   Mon May 2 13:30:52 2022 JCUpdateCoil DriversCoil Drivers Update

The DCC has been updated, along with the modified schematic. Links have been attached.

Quote:

18 (9 pairs) Coil Drivers have been modified. Namely ETMX/ITMX/ITMY/BS/PRM/SRM/MC1/MC2/MC3.

 

ETMX Coil Driver 1 (UL/LL/UR)now has R=100 // 1.2k ~ 92Ohm for CH1/2/3        S2100624                     ETMX Coil Driver 2 (LR/SD)now has R=100 // 1.2k ~ 92Ohm for CH3      S2100631

ITMX Coil Driver 1 (UL/LL/UR)now has R=100 // 1.2k ~ 92Ohm for CH1/2/3         S2100620                      IMTX Coil Driver 2 (LR/SD)now has R=100 // 1.2k ~ 92Ohm for CH3      S2100633

ITMY Coil Driver 1 (UL/LL/UR)now has R=100 // 1.2k ~ 92Ohm for CH1/2/3         S2100623                   ITMY Coil Driver 2 (LR/SD)now has R=100 // 1.2k ~ 92Ohm for CH3      S2100632

BS Coil Driver 1 (UL/LL/UR)now has R=100 // 1.2k ~ 92Ohm for CH1/2/3            S2100625                     BS Coil Driver 2 (LR/SD)now has R=100 // 1.2k ~ 92Ohm for CH3      S2100649

PRM Coil Driver 1 (UL/LL/UR)now has R=100 // 1.2k ~ 92Ohm for CH1/2/3         S2100627                      PRM Coil Driver 2 (LR/SD)now has R=100 // 1.2k ~ 92Ohm for CH3      S2100650

SRM Coil Driver 1 (UL/LL/UR)now has R=100 // 1.2k ~ 92Ohm for CH1/2/3         S2100626                        SRM Coil Driver 2 (LR/SD)now has R=100 // 1.2k ~ 92Ohm for CH3      S2100648

MC1 Coil Driver 1 (UL/LL/UR)now has R=100 // 1.2k ~ 92Ohm for CH1/2/3         S2100628                        MC1 Coil Driver 2 (LR/SD)now has R=100 // 1.2k ~ 92Ohm for CH3      S2100651

MC2 Coil Driver 1 (UL/LL/UR)now has R=100 // 1.2k ~ 92Ohm for CH1/2/3       S2100629                        MC2 Coil Driver 2 (LR/SD)now has R=100 // 1.2k ~ 92Ohm for CH3         S2100652

MC3 Coil Driver 1 (UL/LL/UR)now has R=100 // 1.2k ~ 92Ohm for CH1/2/3        S2100630                        MC3 Coil Driver 2 (LR/SD)now has R=100 // 1.2k ~ 92Ohm for CH3         S2100653

 

 

Will be updating this linking each coil driver to the DCC

 

  16822   Mon May 2 10:00:34 2022 JordanUpdateVACTP3 Forepump tip seal replacement

[JC, Jordan]

Jordan recieved the new tip seal Friday afternoon and I continued the replacement process in the morning. Finishing up, we proceeded to test the pump in the Clean and Bake room. The pump's pressure lowered to 110 mTorr, and we continue pumping so the seal can recieve a good fitting.

Update: We have confirmed the pump is working great and have reinstalled this back into the vacuum system.                                                                                                                   Note: The same O-Rings were used.

Quote:

Jordan, JC

While the pumpspool is vented, I thought it would be a convenient time to change out the tip seal on the TP3 forepump. This one had not been changed since 2018, so as preventative maintence I had JC remove the pump and begin cleaning/installing the new tip seal.

Unfortunately the tip seal broke, but I have ordered another. We should have this pump ready to go late next week. If one is needed sooner, there is a spare IDP 7 pump we can install as the TP3 forepump.

 

  16821   Sat Apr 30 13:29:33 2022 PacoUpdateBHDPOX Beam issues --- fixed, ASL alignment on vertex chamber

[Paco, Anchal, Yuta]

We opened the BSC and ITMX chamber in the morning (Friday) to investigate POX11 beam clipping. We immediately found that the POX11 beam was clipping by the recently installed cable posts, so luckily no major realingment had to be done after reinstalling the cable post in a better location.


Because we had the BSC open, we decided to steer the AS1 mirror to align the AS path from ITMY all the way to the vertex chamber.  Relatively small AS1 offsets (of ~ 2000 counts each) were added on PIT / YAW to center the beam on ASL (there is slight clipping along PIT, potentially because of the AS2 aperture. We then opened the vertex chamber and located the AS beam with relative ease. We decided to work on this chamber, since major changes propagate heavily downstream (simply changing the IMC pointing).

Anchal removed old optics from the vertex chamber and we installed the steering pair of mirrors for AS path. This changed the balance of the vertex table by a lot. By using the MC REFL camera beam spot we managed to coarsely balance the counterweights and recover the nominal IMC injection pointing. Simply reenabling the IMC autolocker gave us high transmission (~ 970 counts out of the typical 1200 these days).

The final IMC alignment was done by Anchal with delicate PIT motion on the input injection IMC miror to maximize the transmission (to our satisfaction, Anchal's motion was fine enough to keep the IMC locked). The end result was quite satisfying, as we recovered ~ 1200 counts of MC transmission.

Finally, we looked at the arm cavity transmission to see if we were lucky enough to see flashing. After not seeing it, we adjusted TT1 / TT2 to correct for any MMTT1 pitch adjustment needed after the vertex table rebalancing. Suprisingly, we didn't take too long and recovered the nominal arm cavity pointing after a little adjustment. We stopped here, but now the vertex table layout is final, and AS beam still needs to be aligned to the vertex in-air table.

  16820   Fri Apr 29 08:34:40 2022 JCUpdateVACRGA Pump Down

Jordan and I, in order to start pumpig down the RGA Volume, we began by opening V7 and VM. Afterwards, we started RP1 and RP3. After this, the pressure in the line between RP1, RP3, and V6 dropped to 3.4 mTorr. Next, we tried to open V6, although an error message popped up. We haven't been able to erase it since. But we were able to turn on TP2 with V4 closed. The pressure in that line is reporting 1.4 mTorr.

 

PRP on the sitemap is giving off an incorrect pressure for the line between RP1, RP3, and V6. This is verified by the pressure by the control screen and the physical controller as well. 

Attachment 1: Screen_Shot_2022-04-29_at_8.46.53_AM.png
Screen_Shot_2022-04-29_at_8.46.53_AM.png
  16819   Thu Apr 28 18:07:04 2022 AnchalUpdateBHDRestoring arm alignment at EOD

Restored arm algiment to get 0.8 max flashing on YARM and 1 max flashing on XARM. I had to move input alignment using TT2-PR3 pair and realign YARM cavity axis using ITMY-ETMY pair.

I would like to advertise this useful tool that I've been using for moving cavity axis or input beam direction. It's a simple code that makes your terminal kind of videogame area. It moves two optics together (either in same direction or opposite direction) by arrow key strokes (left, right for YAW, up, down for PIT). Since it moves two optics together, you actually control the cavity axis or input beam angle or position depening on the mode.

 

  16818   Thu Apr 28 17:45:53 2022 AnchalUpdateBHDPOX Beam issues

[Anchal, Paco]

We investigated the low power issue with POX11 photodiode.

  • We used a power meter to confirm that above 12 uW of power is reaching the ITMX oplev table.
  • But the power reaching the photodiode was only 3 uW, because the 2 steering mirrors used were discarding half of the light. These mirrors were taken from former POP setup and are probably BS or meant for different incidence angle/polarization.
  • We used flashlight to test that the photodiode gives a response, so it is not dead.
  • We also put in a db15 breakout board in line to the PD and confirmed that power input is correct, temperature is nominal, enable is ON, and DC out is responsive to flashlight.
  • So we decided to redo the path with Y1-45P 1inch mirrors. I found two such mirrors in the lab.
  • When setting up the path again, I realized that the beam shape is not even. I took a photo of the beam on the card (see attachment 1) and indeed the POX beam that is coming to the table is half clipped.
  • So tomorrow, we'll open the chamber and find out where this is happening. For now, I've setup a nominal path to the POX11 photodiode, but we'll tune it after we get a clear POX beam on the table.
Attachment 1: PXL_20220429_002703686.jpg
PXL_20220429_002703686.jpg
  16817   Thu Apr 28 11:53:10 2022 AnchalSummaryBHDPOP_SM4 and POP_SM5 Assembly

POP_SM4

I tried out this stack today and found some change of plans.

  • The attachment in elog 40m/16640 says to use 1/4-20 silver coated non-vented screws for joining BA2V to PLS-T238, however PLS-T238's bottom hole is a blind hole and is not vented. So we actually required <1 in long silver plated vented 1/4-20 cap screw for this purpose. Jordan and I were only able to find the correct length silver plated screw but it is not vented. So we decided to make a venting hole on the post from the side.
  • I had to use a 0.14" spacer as washer between PLS-T238 and BA1V. The 1" post shim that Koji got for this purpose had too small a hole in the center to let the 8-32 screw pass (I know, weird). but I think we had 3 spare 0.14" ad we bought 10 when we required 7, so we should be good.
  • The attachment in elog 40m/16640 also says to use 8-32 silver coated vented setscrew for joining TR-1.5 to LMR1V mount. I found one vented silver coated set screw nearby in the clean room but it turned out to be too long. Worse, I overtightened the setscrew when trying this connection which damaged the inner threads of one of the LMR1V. So we need to buy one more LMR1V (maybe an additional spare too) for future installation of OMC1R1/OMC2R1.
    • Then Jordan and I searched for a smaller silver coated and vented 8-32 setscrew but didn't find any. Jordan also noted that LMR1V is an aluminium mount and we should not use silver coated setscrew with it. Since the TR-1.5 mount is ok to be sacrificed if a cold weld happens, we'll just use an uncoated SS 8-32 setscrew to join TR-1.5 and LMR1V. We could not such vented setscrews, but we have plenty of non-vented once. So Jordan is going to make a venthole in TR1.5 top end as well. LMR1V already has a vent hole on its side.

tl,dr; Jordan is preparing PLS-T238 and TR-1.5 with venting holes and C&B and they would be ready by tomorrow. I have collected all other parts for assembly, still looking for the mirror but I know other lab members know where it is, so no big issue there.


POP_SM5

The assemly of this mirror is complete. A slight change here as well, we were supposed to use the former POYM1 (Y1-2037-0) mirror for POP_SM5 but I could not find it. It was stored on the right most edge of the table (see 40m/16450), but it is not there anymore. I found another undocumented mirror on the flow bench on the left edge marked (2010 July: Y1-LW1-2037-UV-0-AR) which means this mirror has a wedge of 1 degree and an AR coating as well. We do not need or care about the wedge or AR coating, so we can use this mirror for POP_SM5. Please let me know if someone was saving this mirror for some other purpose.


I'll finish assembly of POP_SM4 tomorrow and install them in ITMX chamber and resurrect POP path.

Quote:

Here is more detail of the POP_SM4 mount assembly.

It's a combination of BA2V + PLS-T238 + BA1V + TR-1.5 + LMR1V + Mirror: CM254-750-E03
Between BA1V and PLS-T238, we have to do a washer action to fix the post (8-32) with a 1/4-20 slot. Maybe we can use a 1" post shim from thorlabs/newport.
Otherwise, we should be able to fasten the other joints with silver-plated screws we already have/ordered.

I think TR-1.5 (and a shim) has not been given to Jordan for C&B. I'll take a look at these.

 

  16816   Thu Apr 28 09:12:18 2022 AnchalUpdateBHDRestoring arm alignment

This would be a daily first task in the morning. We'll need to check the status of arm alignment and optimize it back to maximum every morning for the rest of the day's work.

Today, when I came, on openin gthe PSL shutter, IMC was aligned good, both arms were flashing but YARM maximum transmission count was around 0.7 (as opposed to 1 from yesterday) and XARM maximum transmission count was 0.5 (as opposed to 1 from yesterday). I did not change the input alignment to the interferometer. I only used ITMY-ETMY to regain flashing count of 1 on YARM and used BS and tehn ITMX-ETMX to regain flashing count of 0.9 to 1 in XARM.

Even thought the oplevs were centered yesterday, I found the oplev had drifted from the center and the optimal position also is different for all ooptics except EMTY and BS. It is worth nothign that in optimal position both PIT and YAW of ITMY and ITMX are off by 70-90 uradians and ETMX Pit oplev is off by 55 uradians.

 

  16815   Wed Apr 27 16:28:57 2022 AnchalSummaryBHDBHD Upgrade - Retreiving arm cavity alignment

[Anchal, Paco, JC]

We had to open ITMY, ETMY chamber doors to get the cavity aligned again. Once we did that, we regained cavity flashing and were able to align the input injection and cavity alignment to get transmission flashing to 1.0 (C1:LSC-TRY_OUT_DQ). JC later centered both ITMY and ETMY oplevs. The ITMY oplev had become completely out of range.

We also opened ITMX, ETMX chamber doors to get Xarm alignment. Again, it seems that ITMX had moved a lot due to cable post installation.

To be continued

  16814   Wed Apr 27 10:05:55 2022 JCUpdateCoil DriversCoil Drivers Update

18 (9 pairs) Coil Drivers have been modified. Namely ETMX/ITMX/ITMY/BS/PRM/SRM/MC1/MC2/MC3.

 

ETMX Coil Driver 1 (UL/LL/UR)now has R=100 // 1.2k ~ 92Ohm for CH1/2/3        S2100624                     ETMX Coil Driver 2 (LR/SD)now has R=100 // 1.2k ~ 92Ohm for CH3      S2100631

ITMX Coil Driver 1 (UL/LL/UR)now has R=100 // 1.2k ~ 92Ohm for CH1/2/3         S2100620                      IMTX Coil Driver 2 (LR/SD)now has R=100 // 1.2k ~ 92Ohm for CH3      S2100633

ITMY Coil Driver 1 (UL/LL/UR)now has R=100 // 1.2k ~ 92Ohm for CH1/2/3         S2100623                   ITMY Coil Driver 2 (LR/SD)now has R=100 // 1.2k ~ 92Ohm for CH3      S2100632

BS Coil Driver 1 (UL/LL/UR)now has R=100 // 1.2k ~ 92Ohm for CH1/2/3            S2100625                     BS Coil Driver 2 (LR/SD)now has R=100 // 1.2k ~ 92Ohm for CH3      S2100649

PRM Coil Driver 1 (UL/LL/UR)now has R=100 // 1.2k ~ 92Ohm for CH1/2/3         S2100627                      PRM Coil Driver 2 (LR/SD)now has R=100 // 1.2k ~ 92Ohm for CH3      S2100650

SRM Coil Driver 1 (UL/LL/UR)now has R=100 // 1.2k ~ 92Ohm for CH1/2/3         S2100626                        SRM Coil Driver 2 (LR/SD)now has R=100 // 1.2k ~ 92Ohm for CH3      S2100648

MC1 Coil Driver 1 (UL/LL/UR)now has R=100 // 1.2k ~ 92Ohm for CH1/2/3         S2100628                        MC1 Coil Driver 2 (LR/SD)now has R=100 // 1.2k ~ 92Ohm for CH3      S2100651

MC2 Coil Driver 1 (UL/LL/UR)now has R=100 // 1.2k ~ 92Ohm for CH1/2/3       S2100629                        MC2 Coil Driver 2 (LR/SD)now has R=100 // 1.2k ~ 92Ohm for CH3         S2100652

MC3 Coil Driver 1 (UL/LL/UR)now has R=100 // 1.2k ~ 92Ohm for CH1/2/3        S2100630                        MC3 Coil Driver 2 (LR/SD)now has R=100 // 1.2k ~ 92Ohm for CH3         S2100653

 

 

Will be updating this linking each coil driver to the DCC

  16813   Tue Apr 26 16:23:22 2022 TommyUpdateElectronicsRFSoC2x2 MTS

We connected a 8 MHz signal generator to the device in order to sync up the ADCs and DACs and hopefully get phase data. 

Some things to note:

  • RF Manual (143)- Need to use XRFDC SYSREF for update event
  • RF Manual (171)- Synchronization steps require us to first enable all clocks and sysref generators (via xrfdc package)
  • RF Manual (173)- Sysref requirments, not clear if PL is syncing as needed.
  • RF Manual (181)- XRFDC example code, see also https://github.com/Xilinx/embeddedsw/blob/master/XilinxProcessorIPLib/drivers/rfdc/examples/xrfdc_mts_example.c

Xilinx RF Manual: https://docs.xilinx.com/v/u/2.4-English/pg269-rf-data-converter

  16812   Mon Apr 25 18:00:03 2022 Ian MacMillanUpdateUpgradeCable supports update

I have designed new cable supports for the new ribbon cables running up the side of the tables in the vacuum chambers. 

The clamps that I have designed (shown in basic sketch attachment 1) will secure the cable at each of the currently used cable supports. 

The support consists of a backplate and a frontplate. The backplate is secured to the leg of the table using a threaded screw. The frontplate clamps the cable to the backplate using two screws: one on either side. Between two fascinating points, the cable should have some slack. This should keep the cable from being stiff and help reduce the transfer of seismic noise to the table. 

It is possible to stack multiple cables in one of these fasteners. Either you can put two cables together and clamp them down with one faceplate or you can stack multiple faceplates with one cable between each faceplate. in this case the stack would go backplate then cable then faceplate then cable then the second faceplate. this configuration would require longer screws.

The exact specifics about which size screws and which size plates to use still have not been measured by me. But it will happen

Attachment 1: Chamber_Leg_Ribbon_Cable_Attachments.pdf
Chamber_Leg_Ribbon_Cable_Attachments.pdf
  16811   Mon Apr 25 17:24:06 2022 AnchalUpdateCDSDAQ still down

I investigated this issue today. At first, it seemed that only new suspension testpoints are inaccessible. I was able to use diaggui for a measurement on MC2. The DAQ network cable between 1X4 and 1Y1 was tied and is very taught now (we should relieve this as soon as possible, best solution is to lay down a longer cable over the bridge). My hypothesis is that the DAQ network might have broken while tying this cable and it probably did not come back since then.

The simplest solution would have been to restart c1su2 models. As I restarted those models though, I found that c1lsc and c1sus models failed. This is very unusual as c1su2 models are independent and share nothing with the other vertex models. So I had to restart all the FE computers eventually. But this did not solve the issue. Worse, now the DAQ isn't working for the vertex machiens as well.

Next step was to try restarting fb1 itself. We switched off all the FE computers, restarted fb1, stopped daqd_* processes, reloaded gpstime module, restarted open-mx, mx, nds and daqd_* process. But the mx.service failed to load with following error message:

● mx.service - LSB: starts MX driver for Myrinet card
   Loaded: loaded (/etc/init.d/mx)
   Active: failed (Result: exit-code) since Mon 2022-04-25 17:18:02 PDT; 1s ago
  Process: 4261 ExecStart=/etc/init.d/mx start (code=exited, status=1/FAILURE)

Apr 25 17:18:02 fb1 mx[4261]: Loading mx driver
Apr 25 17:18:02 fb1 mx[4261]: insmod: ERROR: could not insert module /opt/mx/sbin/mx_mcp.ko: File exists
Apr 25 17:18:02 fb1 mx[4261]: insmod: ERROR: could not insert module /opt/mx/sbin/mx_driver.ko: File exists
Apr 25 17:18:02 fb1 systemd[1]: mx.service: control process exited, code=exited status=1
Apr 25 17:18:02 fb1 systemd[1]: Failed to start LSB: starts MX driver for Myrinet card.
Apr 25 17:18:02 fb1 systemd[1]: Unit mx.service entered failed state.

(Ignore the timestamp above, I ran the command again to capture the error message.)

However, I was able to start all the FE models without any errors and daqd processes are also all running without showing any errors. Everything is green in CDS screen with no error messages. But the only thing still wrong is mx.service which is not running.

From my limited knowledge and experience, mx.service is a one-time script that mounts mx devices in /dev and loads the mx driver. I tried running the script /opt/mx/sbin/mx_start_stop :

controls@fb1:/opt/mx/sbin 1$ sudo ./mx_start_stop start
Loading mx driver
insmod: ERROR: could not insert module /opt/mx/sbin/mx_mcp.ko: File exists
insmod: ERROR: could not insert module /opt/mx/sbin/mx_driver.ko: File exists

This gave the same error. On searching little bit online, "insmod: ERROR; cound not insert module" error comes up when the kernel version of the driver doesnot match the Linux kernel (whatever that means!). Such deep issues should not appear out of nowhere in a previosuly perfectly runnig system. I'll check around more what changed in fb1, network cables etc.

  16810   Mon Apr 25 16:57:57 2022 AnchalUpdateBHDCable post installation

[Anchal, Tega, JC]

We installed cable posts in ITMY, BS, and ITMX chambers for all the new suspensions. Now, there is no point where the OSEM connections are hanging freely.

In BS chamber, we installed one post for LO2 near the north edge of the table and another post for PR3 on the Western edge with the blue cable running around the table on the floor.

In ITMY chamber, we installed the cable post in between AS1 and AS4 with the blue cables running around the table on the floor. This is to ensure the useful part of the table remains empty for future and none of the OSEM cables are taught in air.

 

  16809   Mon Apr 25 14:49:02 2022 KojiUpdateGeneralNitrogen Tank

For your (and mine) info:

N2 pressure can be monitored on the 40m summary page: https://nodus.ligo.caltech.edu:30889/detcharsummary/day/20220425/vacuum/
(you need to hit "today" to go to the current status)

 

ELOG V3.1.3-